Christiana's birthday is four more than three times Rolando's birthday. Monique's birthday is one more than three times Rolando's. If they all add up to 54, what are their birthdays?
Answer by jonvaliente(64)

Let x=Rolando's birthday
3x+4=Christina's birthday
3x+1=Monique's birthday
x+(3x+4)+(3x+1)=54
7x+5=54
7x+5-5=54-5
7x=49
x=7
Rolando's birthday is on the 7th, Christina's is on the 25th and Monique's is on the 22nd